# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v23

@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [73]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [563] where they used, the fasttext model trained by Facebook [116, 117, 118], on a RASA126126126https://rasa.com/ [564] chat bot. A Sinhala chat bot for train information was proposed by Harshani [565]. Similarly, the tool proposed by Chandrasena et al. [566] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [567] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [568] used off the shelf RASA NLU Engine [564] and Microsoft Bot Network [569] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [570]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [571] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [375] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
@@ 3.23 Sinhala Sign Language
The work by Samarasinghe et al. [582] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [607] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [586] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [584] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [608] and Herath and Ishanka [609] integrated MediaPipe [610] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [608] only used LSTM while Herath and Ishanka [609] used LSTM, CNN, and CNN-LSTM. Rishan et al. [579] uses Leap Motion Controller127127127https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [583] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [621] and later by Smith [622] using an XGB classifier and a CRF model building on their previous work [623], which analysed such data. Shanmugalingam and Sumathipala [624] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [485]. They later used it for NMT in Sinhala-English Code-Mixed text [484]. As discussed in Section 3.10Rathnayake et al. [232] used adapter-based [233, 234, 235, 236, 237, 238] fine-tuning on XLM-R [239], for classifying code-mixed and code-switched Sinhala text. Hettigoda [242] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [262] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [244] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [570] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [420] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [625] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by Udawatta et al. [626] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection. Uthpala and Thirukumaran [264] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [627] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [628] assists children to learn to speak without stuttering. Nethmi et al. [629] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [630] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [631] uses SVM and LSTm to identify phonological disorders from Sinhala speech.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [632]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [104] has published two Sinhala BERT models: SinBERT-large128128128https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small129129129https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo130130130https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [633] on the OSCAR [148, 149] dataset.
Üstün et al. [634] introduced Aya, a massively multilingual generative language model131131131https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [152].
@@
Wickramasinghe and de Silva [141] used Sinhala-English dictionary data [140] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub132132132https://bit.ly/3t3SKu7.
@@
Jayakody and Dias [635] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude133133133https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o134134134https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)135135135https://llama.meta.com/docs/get-started/ and Mistral 7B136136136https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2137137137https://falconllm.tii.ae/ and Gemini138138138https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala.
@@ 3.29 Text Simplification
The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [636] included the NSINA [105] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data139139139https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data140140140https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [637] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [638] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [639] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [640] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
@@ 3.30 Miscellaneous Applications
In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them.
Dissanayake and Hettige [641] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [642] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha. Romero et al. [154] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data141141141https://huggingface.co/datasets/afaji/cvqa and leader board142142142https://eval.ai/web/challenges/challenge-page/2305/overview. On the other hand, Amarasinghe and Ranathunga [643] used an ontology-based approach to generate Sinhala essay questions. The study by Kao and Ilmini [644] used LSTM to generate Sinhala lyrics. Bandara et al. [645] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala. However, they did not train a GTP-2 model in Sinhala. Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API. The work by Liyanage and Ranathunga [646, 647] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages. The follow-up work by Niyarepola et al. [648] used pre-trained mBART [491] and mT5 [310, 649] models for the same task and showed better results.
@@
Ranasinghe et al. [656] trained a Sinhala image caption generator model on Google translated MS COCO [657] dataset and tested it on Google translated Flickr8K [658] and MS COCO datasets.
@@
Kumari and Hettiarachchi [660] proposed an algorithm for Sinhala topic modelling based on LDA [241] and RAKE [661]Batawalaarachchi [662] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [663] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [664] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [665] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [666] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [667] have created a small Sinhala treebank of 100 sentences. Burchell et al. [139] show that they obtain better results for the LID task for most of the languages compared to NLLB [130]. However, for Sinhala the results by Burchell et al. [139] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Minixhofer et al. [668] have used Sinhala from OPUS100 [127] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. The study by Petrov et al. [669] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [649] and mT5 [310, 649] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [670]. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [671]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [672]. Senarathne [673] introduced a model to translate Sinhala pseudocode to computer code written in C#. Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [625] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [674] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [675] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. The study by [208] used NER to cluster Sinhala news articles.

### v24

@@ 3.21 Chat Bots
A simple Sinhala chat bot which utilizes a small knowledge base has been proposed by Hettige and Karunananda [78]. A study on the effect of word embeddings on a Sinhala chatbot was conducted by Gamage et al. [604] where they used, the fasttext model trained by Facebook [120, 121, 122], on a RASA148148148https://rasa.com/ [605] chat bot. A Sinhala chat bot for train information was proposed by Harshani [606]. Similarly, the tool proposed by Chandrasena et al. [607] serves as a chat bot-based recommendation system for Sri Lankan traditional dancers. The chat bot discussed by Kumanayake [608] has the very specific purpose of answering user inquiries about the degree programs at University of Ruhuna. Avishka et al. [609] used off the shelf RASA NLU Engine [605] and Microsoft Bot Network [610] to set up a generic Sinhala chat bot architecture. They demonstrated the effectiveness of their architecture by creating a food ordering chat bot. A web-based code-less chat bot development platform for Sinhala was proposed by Dissanayake et al. [611]. Further, they claimed that their system can handle OOV tokens as well as Sinhala-English code-switching. The work by Dasanayaka and Warnajith [612] used a deep learning Intent Mapping (IM) model to map consumer responses in their Sinhala chat bot framework. Rajapakshe et al. [405] proposed a Sinhala conversational interface for the purpose of scheduling medical appointments and giving medical advice. The chat bot component was in the back end while the front end contained speech recognition and voice synthesizer components.
@@ 3.23 Sinhala Sign Language
The work by Samarasinghe et al. [623] used a CNN model to convert Sinhala sign language to Sinhala text. Kumar et al. [648] introduced EasyTalk a tool that translates Sri Lankan sign language to text as well as converts English text into Sri Lankan Sign Language. The work by dha [627] used CNN to classify features of static Sinhala signs while dynamic signs were modelled as a tree structure of static signs. Dahanayaka et al. [625] also used CNN in their multi-model approach to recognize Sinhala sign language. Both the approaches by Perera and Kumarika [649] and Herath and Ishanka [650] integrated MediaPipe [651] with Deep Neural Networks to interpret Sri Lankan sign language. Perera and Kumarika [649] only used LSTM while Herath and Ishanka [650] used LSTM, CNN, and CNN-LSTM. Rishan et al. [620] uses Leap Motion Controller149149149https://leap2.ultraleap.com/leap-motion-controller-2/ to translate Sri Lankan sign language to Sinhala text. Dissanayake et al. [624] have introduced Utalk, a mobile app to interpret Sri Lankan Sign Language.
@@ 3.26 Sinhala-English Code-Mixing
The problem of recognizing Sinhala and English code-mixed data where the Sinhala text is written in Singlish was explored by Smith and Thayasivam [662] and later by Smith [663] using an XGB classifier and a CRF model building on their previous work [664], which analysed such data. Shanmugalingam and Sumathipala [665] also attempted to identify the language in Sinhala-English code-mixed text using Support Vector Machines (SVM), Naive Bayes, Logistic Regression, Random Forest, and Decision Trees. A dictionary based approach to standardize Sinhala Code-Mixed text was proposed by Kugathasan and Sumathipala [518]. They later used it for NMT in Sinhala-English Code-Mixed text [517]. Shakir and Deuber [666] claims to have developed a corpus of South Asian languages in the context of code-mixing which includes Sinhala. However, this is not publicly available. It should further be noted that they report Sinhala data using the heading Sri Lanka rather than Sinhala. In their follow-up work Shakir and Deuber [667] claim to extend this code-mixed data set to include meme data.
As discussed in Section 3.10Rathnayake et al. [253] used adapter-based [254, 255, 256, 257, 258, 259] fine-tuning on XLM-R [260], for classifying code-mixed and code-switched Sinhala text. Hettigoda [263] classified English-Sinhala code-mixed comments from Facebook. As discussed in Section 3.11Aththanayaka and Naleer [283] utilized traditional machine learning methods for sentiment analysis on Sinhala-English code-mixed text from social media. Chathuranga and Ranathunga [265] used capsule-based methods to classify Sinhala-English code-mixed data.  Dissanayake et al. [611] claim that the web-based code-less chatbot development platform for Sinhala proposed by them is capable of handling Sinhala-English code-switching. The feature set derived by Walawage and Ranathunga [451] for text on social media images (memes) attempts to separately identify Sinhala and English text. The study by Fazal and Farook [668] used classical machine learning algorithms on TF-IDF features to predict depression in Sinhala-English code-mixed data from Twitter and Facebook. The study by Udawatta et al. [669] shows that prompt-based learning of pre-trained language models (PLMs) outperforms full fine-tuning of PLMs on Code-mixing and code-switching (CMCS) English-Sinhala data across various NLP tasks such as sentiment classification, hate-speech detection, and humour detection. Uthpala and Thirukumaran [285] claims to have created a sentiment-annotated Sinhala-English code mixed dataset using comments from YouTube videos. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset.
@@ 3.27 Language Acquisition and Evaluation
Sandathara et al. [670] proposed a system which they named Arunalu that they claimed to use Voice recognition, Natural Language Processing, Machine Learning, and Deep Learning concepts to help individuals with dyslexia overcome problems of reading Sinhala. The learning bot MiMi proposed by Vithana et al. [671] assists children to learn to speak without stuttering. Nethmi et al. [672] introduced Narrataa, a mobile application targeted at developing the Sinhala vocabulary of children by using computer vision to caption images in Sinhala. The mobile application developed by Saranga et al. [673] helps individuals learn Sinhala letters with the help of images, colours, and shapes. The application also attempts to generate stories in Sinhala. The Katha App developed by Wijesooriya et al. [674] uses SVM and LSTm to identify phonological disorders from Sinhala speech. Francis et al. [675] claims to use Sinhala OCR and PoS tagging in their effort to build the SinhaLearn, a system for improving Sinhala Proficiency of grade 5 scholarship students. Praveen et al. [362] also claims to have created a web-based education system that helps grade 5 scholarship students who are learning the Sinhala language. They claim that they have created a Sinhala Spell checker with 82% accuracy and a Sinhala TTS system with 98% accuracy within this system.
@@ 3.28 Embeddings, LLMs, and Other Language Models
A two-tiered model to embed Sinhala sentences was proposed by Weeraprameshwara et al. [676]. According to their results, Seq2Seq GRU with attention run on fastText word embedding obtains the best results for Sinhala sentence embedding. Dhananjaya et al. [107] has published two Sinhala BERT models: SinBERT-large150150150https://huggingface.co/NLPC-UOM/SinBERT-large and SinBERT-small151151151https://huggingface.co/NLPC-UOM/SinBERT-small. A smaller BERT model named SinhalaBERTo152152152https://huggingface.co/keshan/SinhalaBERTo has also been published by a different individual but does not have an accompanying research paper. This BERT model has been created by training vanilla RoBERTa [677] on the OSCAR [155, 156] dataset. Demotte et al. [272] have released the word embeddings153153153https://drive.google.com/drive/folders/1CgqPIloW5DKLVj3WViMitIhtrswOv_qR as well as the code154154154https://github.com/theisuru/sentiment-tagger for their word embedding-based Sinhala sentiment analyzer. Gurgurov et al. [678] included Sinhala as one of their 87 low-resource languages for which they provided word embeddings in their LowREm repository. They credit Demotte et al. [272] as their source for Sinhala embeddings.
@@
Wickramasinghe and de Silva [148] used Sinhala-English dictionary data [147] for obtaining Sinhala-English embedding alignment. Further, they established benchmarks for the Sinhala-English embedding alignment task and compared the results against other language pairs such as English-Spanish, English-French, English-German, English-Russian, and English-Chinese. The relevant code and data can be accessed on GitHub155155155https://bit.ly/3t3SKu7. Subsequent work by Rathnayake et al. [679] also attempts BLI for Sinhala English but uses UVecMap (Unsupervised VecMap [680]) as the baseline. They report that UVecMap combined with Linear Transformation archives the best result of 33.18 for En-Si Word2Vec embedding, while UVecMap combined with Linear Transformation and CSCBLI [681] archives the best result of 29.49 for En-Si FastText embedding.
@@
Üstün et al. [171] introduced Aya, a massively multilingual generative language model156156156https://huggingface.co/CohereForAI/aya-101 trained on their Aya data set [159]. Jayakody and Dias [682] conducted a comparative analysis on LLMs for Sinhala. In it, they observed that the small (7̃B) models of Claude157157157https://docs.anthropic.com/en/docs/about-claude/models and GPT-4o158158158https://openai.com/index/hello-gpt-4o/ handled Sinhala better than Llama 3 (8B)159159159https://llama.meta.com/docs/get-started/ and Mistral 7B160160160https://docs.mistral.ai/. Further, they reported that they had to exclude Falcon 2161161161https://falconllm.tii.ae/ and Gemini162162162https://deepmind.google/technologies/gemini/ from the study due to poor or non-existent support for Sinhala. Chang et al. [683] introduced Goldfish, a collection of monolingual language models163163163https://huggingface.co/goldfish-models which includes Sinhala trained on data from MADLAD [160], NLLB [134], and Glot500 [109]. The work by Ayesha et al. [684] used a fined-tuned Llama 3.1 (8B) model [685] to predict student performance by analyzing textual entries created by students in Sinhala and English.
Vayani et al. [164] analyzed the performance of a number of LLMs over 100 languages that included Sinhala. The basic dataset was created by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165]. This was further extended with culturally specific data points by searching for images under country name, language name, and cultural category. This dataset, which they named ALM-Bench164164164https://huggingface.co/datasets/MBZUAI/ALM-Bench, is publicly available. They made a number of observations in this aspect. First, they observed that for the Q&A task, the performance of GPT-4o [166] for Sinhala is better than Qwen2-VL [686]. Next, they report that including images in the question further improved Sinhala results of GPT-4o by 38.7%. In the error analysis of the same, they report that 36% of the errors are language errors (fluency), 31% of the errors are due to lack of knowledge, and 18% of the errors are due to a lack of cultural understanding.
@@ 3.29 Text Simplification
The text simplification Shared task of Building Educational Applications (BEA) at MLSP 2024 [687] included the NSINA [108] Sinhala corpus. The aggregated data from the shared task is available online as Train-Test data165165165https://github.com/MLSP2024/MLSP_Data/ and cleaned gold data166166166https://huggingface.co/datasets/MLSP2024/MLSP2024. Enomoto et al. [688] experimented with GPT-4 and noted that re-ranking systems outperformed the base system for Sinhala while the opposite was true for all the other nine languages. Goswami et al. [689] also tested GPT-4 (Terbo) for this task against the top 10 suggestion model and reported better results for GPT-4 for Sinhala than the top 10 suggestion model. Seneviratne and Suominen [690] compared multiple models for the English-Sinhala language pair simplification for the lexical complexity prediction task and lexical substitution task. In the submission by Cristea and Nisioi [691] word frequency and other meta-knowledge are not used for Sinhala text simplification while they are used for the other languages in the task. Their Sihala models are trained only using Sinhala text Google translated to English. As such the results they report for Sinhala for this task is quite weak compared to their results for the other languages.
3.30 Tokenizing
The study by Petrov et al. [692] compared the performance 17 off-the-shelf language tokenizers on a large number of languages including Sinhala. They have found that some languages such as English have shorter tokenizing lengths (close to $1.0$) across the board while some can raise to undesirable heights (they observe $15.0$ as the highest). They report undesirable numbers for Sinhala, mostly in the $8.83$ to $12.86$ range. They report reasonable results on MBart50 [693] and mT5 [336, 693] with $1.35$ and $1.66$ respectively. The best result ($1.00$), is given by CANINE [694]. The work by Velayuthan and Sarveswaran [695] showed that the algorithms used in pre-tokenization are more important than the tokenization algorithm itself in the context of LLMs for languages such as Sinhala. Ranasinghe [163] claims to have created a comprehensive tokenizer for Sinhala and have made it publicly available167167167https://github.com/Buddhilive/sltk. However, the tokenizer is heavily dependent on a manually created dictionary to handle compound words. While the authors have made the dictionary publicly available168168168https://github.com/Buddhilive/sltk/blob/main/sltkpy/shared/dict.json, it only contains seven entries. So while it may be adequate to handle Sri Lanka-centric news title data set169169169https://github.com/Buddhilive/sltk/blob/main/data/sin.csv provided with the code, it may not work well in the general domain.
3.31 Detection or Management of Mental Illnesses
Using classical machine learning algorithms on TF-IDF features, Fazal and Farook [668] predicted depression in Sinhala-English code-mixed data from Twitter and Facebook. Hewapathirana and Sumanathilaka [696] introduced EmoScan, a method to detect depression in Romanized Sinhala (Singlish) tweets. The study by Herath and Wijayasiriwardhane [697] used classical machine learning models to predict suicidal ideation from Sinhala Facebook posts. Fazal and Farook [698] compared multiple algorithms to detect depression in Sinhala English code mixed social media posts and reported that ExtraTreesClassifier [699] obtains the best results. Welarathna et al. [413] used CNNs to classify emotions of Sinhala speech by Autistic children. They handle the following emotions: sad, disgust, surprise, neutral, happy, calm, fear, and angry. Jayawardena et al. [700] used a NAO Robot [701] to create an interactive system with Sinhala language content detection (Speech to Text), dialogue management, and voice synthesis (Text-to-Speech) for Autism Intervention in children.
3.32 Sinhala Text Generation
The study by Kao and Ilmini [702] used LSTM to generate Sinhala lyrics. Bandara et al. [703] introduced Sibil AI, a GTP-2 based Children’s story generator for Sinhala. However, they did not train a GTP-2 model in Sinhala. Instead, they trained an English GTP-2 model and generated stories in English which were then translated to Sinhala using Google translate API. Amarasinghe and Ranathunga [704] used an ontology-based approach to generate Sinhala essay questions. Ranasinghe et al. [705] trained a Sinhala image caption generator model on Google translated MS COCO [706] dataset and tested it on Google translated Flickr8K [707] and MS COCO datasets. The work by Liyanage and Ranathunga [708, 709] attempt to use LSTMs for mathematical word problem generation in Sinhala and other languages. The follow-up work by Niyarepola et al. [710] used pre-trained mBART [524] and mT5 [336, 693] models for the same task and showed better results.
@@ 3.33 Miscellaneous Applications
In this section, we discuss NLP tools and research which are either hard to categorize under the above sections or reasonably involve multiples of them. Dissanayake and Hettige [711] implemented a question and answer generator for Sinhala with the limited PoS: pronouns, adjectives, verbs, and adverbs. Jayakody et al. [712] uses simple KNN and SVM methods on a PoS tagged Sinhala corpus to create a question-answering system which they name Mahoshadha. Romero et al. [161] introduced CVQA, a multilingual visual Question Answering benchmark which includes Sinhala. They have publicly released both their data170170170https://huggingface.co/datasets/afaji/cvqa and leader board171171171https://eval.ai/web/challenges/challenge-page/2305/overview. Vayani et al. [164] created a dataset ALM-Bench172172172https://huggingface.co/datasets/MBZUAI/ALM-Bench by machine translating English QA pairs from LLaVA-Bench (In-the-Wild) dataset [165] and extended it by searching for images under country name, language name, and cultural category on the web. They then reported that for Sinhala Q&A (especially visual Q&A), GPT-4o [166] is superior to other tested LLMs. Wansekara and Jayasekara [713] developed a visual Sinhala question and answer system using YOLO for object identification and a BERT model trained on a Sinhala corpus for Sinhala text.
@@
Kumari and Hettiarachchi [720] proposed an algorithm for Sinhala topic modelling based on LDA [262] and RAKE [721]Batawalaarachchi [722] proposed two methods using statistical features to select words to be included in Sinhala document titles. The study by pal [723] conducted a comparative analysis of LDA, LSA, and NMF for Sinhala and concluded that NMF archives best results. Arambewela et al. [724] proposed a Sinhala writing assistant tool utilising CNNS. The study by Bandara et al. [725] has introduced a Sinhala virtual assistant with rule-based decision-making for a pre-set collection of user queries and a general Transformer-based model for general user questions.  Jayaweera et al. [726] used classical machine learning methods to propose dynamic stop word removal. They claim to have released a corpus of 100,000+ Sinhala documents in their paper. But they provide no information on where to obtain this corpus. Liyanage et al. [727] have created a small Sinhala treebank of 100 sentences. Burchell et al. [146] show that they obtain better results for the LID task for most of the languages compared to NLLB [134]. However, for Sinhala the results by Burchell et al. [146] are the same as that of NLLB. Dunn and Edwards-Brown [8] experiment with improving the LID task with the help of a geographical model. However, they conclude that adding the geographical model does not help in the Sinhala LID task given that the LID model and the geographical model already agree with each other 99%. Minixhofer et al. [728] have used Sinhala from OPUS100 [131] dataset to pre-train their multilingual punctuation-agnostic sentence segmentation model. Given that Sinhala is not one of their main target languages, they do not report sentence segmentation results for Sinhala in much detail. However, they do report that their method improves Sinhala sentence splitting F1 score to 86.0 above their SpaCy baseline of 75.8. A meta-study on how language data from social media are used in research was conducted by Hewapathirana [729]. A Sinhala programming language named Helaa based on Java was proposed by Yasasri and Karunarathna [730]. Senarathne [731] introduced a model to translate Sinhala pseudocode to computer code written in C#. The work by De Silva and Athukorala [732] claims to be able to provide Sinhala language explanations to JAVA code as well as generate JAVA code snippets from Sinhala descriptions.
The study by [228] performs best in this regard. used NER to cluster Sinhala news articles. Subasinghe and Samarawickrama [733] analysed the Noto Sans Sinhala173173173https://fonts.google.com/noto/specimen/Noto+Sans+Sinhala font in the context of legibility on small-scale digital device screens. Zhao et al. [734, 735] discussed a Sinhala LLM-based system they utilized to conduct participatory research in Sri Lanka. While they discussed how the LLM system helped their goals and briefly touched on how they utilised a hybrid translation approach of NMT and rule-based methods to circumvent the lack of training data, the specifics of their LLM system are not given.